# Learning to Generate Unit test via Adversarial Reinforcement Learning

### v2

@@ Abstract
Unit testing is a core practice in programming, enabling systematic evaluation of programs produced by human developers or large language models (LLMs). Given the challenges in writing comprehensive unit tests, LLMs have been employed to automate unit test generation, yet methods for training LLMs to produce high-quality unit tests remain underexplored. In this work, we propose UTRL, a novel reinforcement learning (RL) framework that trains an LLM to generate high-quality unit test given a programming instruction. Our key idea is to iteratively train two LLMs, the unit test generator and the code generator, in an adversarial manner via RL: (1) the unit test generator is trained to maximize a discrimination reward, encouraging it to produce tests that reveal faults in the code generator’s solutions; and (2) the code generator is trained to maximize a code reward, encouraging it to produce solutions that pass the unit tests generated by the unit test generator. In our experiment, we demonstrate that unit tests generated by Qwen3-4B trained via UTRL show higher quality compared to unit tests generated by the same model trained via supervised fine-tuning on ground-truth unit tests, yielding code evaluations that more closely align with those induced by the ground-truth tests. Moreover, Qwen3-4B trained with UTRL outperforms frontier models like GPT-4.1 and GPT-4o in generating high-quality unit tests, highlighting the effectiveness of UTRL in training LLMs for the unit test generation.
@@ 2 Related Work
Unit testing has been extensively explored for reliable software development (Fraser & Arcuri, 2011; Alagarsamy et al., 2024). As LLMs show strong capability in programming, recent works have explored the automation of unit test generation using large language models (LLMs), with a focus on evaluating the test generation capability of LLMs (Jain et al., 2024; Wang et al., 2024), iteratively refining the unit test by leveraging coverage analysis or mutation testing as a contextual feedback (Alshahwan et al., 2024; Altmayer Pizzorno & Berger, 2025; Dakhel et al., 2024; Foster et al., 2025), or utilizing co-evolutionary algorithm (Li et al., 2025). Another line of work investigates training LLMs for unit test generation via supervised learning, with a focus on constructing high-quality unit test annotations at scale. CodeRM (Ma et al., 2025) leverages stronger teacher models to generate unit tests automatically, while UTGEN (Prasad et al., 2025) proposes a data collection strategy that perturbs code solutions to create faulty variants, and then collects high-quality test cases that discriminate between the correct and faulty code. The work most closely related to ours is the concurrent study CURE (Wang et al., 2025a), which explores the co-evolution of code generation and test case generation capabilities via reinforcement learning. However, our approach differs from CURE in two key aspects: (1) Unlike CURE, which assumes access to a dataset annotated with ground-truth unit tests, UTRL requires only instruction-code pair dataset, which is readily available at large scale (Li et al., 2023; Lambert et al., 2024; Prasad et al., 2025), and (2) UTRL is a framework aimed at training LLMs to generate an optimal set of test cases, rather than a single test case.
@@ 3.2 Training Unit Test Generator
Figure 3: Overview of the process of computing discrimination reward with respect to unit test $\mathcal{T}$. First, among the $5$ test cases in the unit test $\mathcal{T}$, test cases that pass under the ground-truth code $C^{*}$ (i.e., $T_{1},T_{3},T_{5}$) are filtered, forming a set of functionally valid test cases. Second, the discrimination reward is defined as a ratio of sampled code solutions that do not pass at least one valid test case. In this figure, among 6 sampled code solutions, 4 code solutions ($C_{1},C_{3},C_{4},C_{5}$) do not pass at least one valid test case, resulting in the discrimination reward of 0.667 ($\frac{4}{6}$).
@@ 4.1 Experimental Setup
  • LLM baselines without fine-tuning: We use 4 open-source LLMs (Qwen3-4B, Qwen3-8B, Qwen-14B, Deepseek-Coder-V2-Lite; Qwen et al. 2025; Zhu et al. 2024) and 2 closed LLMs (GPT-4o, GPT-4.1; Hurst et al. 2024; OpenAI 2025) as baselines for unit test generation.
  • Supervised fine-tuning (SFT): We fine-tune Qwen3-4B with instruction–unit test pairs through supervised learning. For this purpose, we use high-quality unit tests from the TACO training dataset $\mathcal{D}_{\text{UT}}$. To investigate whether reasoning signals can improve unit test generation, we construct a reasoning-augmented dataset $\mathcal{D}_{\text{reason+UT}}$ containing unit tests paired with their corresponding reasoning. Specifically, we first generate both reasoning and unit tests for each training task using Gemini-2.5-flash (Google Deepmind, 2025), capable LLMs in code generation. [^1] Instead of attaching reasoning to the original unit tests in TACO, which led to frequent hallucinations and inconsistencies, we regenerate both reasoning and unit tests together. We then filter out functionally invalid test cases and retain the valid reasoning–unit test pairs as target labels for SFT training.
@@
  • Ground-truth (GT) unit tests: We also compare against the unit tests provided in the TACO evaluation dataset (Li et al., 2023), which we consider as ground truth (GT) and regard as an upper bound since they are rigorously verified and consist of a large number of high-quality test cases.
@@
For evaluation, we utilize rigorously verified 945 competitive programming tasks provided in the TACO evaluation dataset (Li et al., 2023), which were collected from various online judge platforms (CodeForces, CodeChef, HackerRank, and HackerEarth). To measure the quality of the generated unit tests, we propose the following evaluation schemes:
  • Best-of-NN improvement: To examine whether the generated unit tests include high-quality test cases, we perform best-of-$N$ sampling for code generation, using the generated unit tests as evaluators. Specifically, we generate $N$ candidate solutions from code LLMs and select the one that passes the largest number of generated unit tests. To evaluate the quality of the selected code, we then measure two metrics: (1) code score, defined as the fraction of test cases in a GT unit test (i,e., high-quality unit tests from the TACO evaluation dataset) that the code solution passes, and (2) code accuracy, which indicates whether the solution passes all test cases in the GT unit test.
@@
We provide more details about the evaluation in Appendix A.4.
@@
Qwen3-4B + UTRL (ours) 0.558 +0.212 14.9 +6.9 0.588 +0.188 17.0 +6.3
Qwen3-14B + UTRL (ours) 0.562 +0.216 15.7 +7.7 0.594 +0.194 17.1 +6.4
GT (upperbound) 0.650 +0.304 21.0 +13.1 0.680 +0.278 24.4 +13.7
@@
4.2 Quality of Unit Tests
Best-of-NN improvement
Table 1 reports the best-of-$N$ improvement achieved by Qwen3-4B and Qwen3-14B trained with UTRL, compared against several baselines. When used as code evaluators for best-of-32 sampling with Qwen3-8B and Qwen3-14B code generators, unit tests produced by Qwen3-4B trained with UTRL yield code accuracies of 14.9% and 17.0%, respectively. For comparison, Qwen3-4B trained with SFT achieves only 11.7% and 14.0%. Moreover, Qwen3-4B trained with UTRL outperforms frontier proprietary models such as GPT-4.1 and GPT-4o, underscoring the effectiveness of UTRL in training LLMs to generate high-quality unit tests. With increased model capacity, Qwen3-14B trained with UTRL attains stronger performance, reaching code accuracies of 15.7% and 17.1%. Notably, UTRL is effective even when paired with closed-source code generators such as GPT-4o (see Table 3 in Appendix A.2 for supporting results).
@@ 4.2 Quality of Unit Tests
Figure 4 presents the fidelity of unit tests generated by Qwen3-4B trained with UTRL, compared against several baselines. Qwen3-14B and Qwen3-4B trained with UTRL achieves a Spearman’s correlation of 0.764 and 0.741, even outperforming GPT-4.1, demonstrating the effectiveness of UTRL in training LLMs to generate unit tests that induce code evaluations consistent with the GT unit tests. Interestingly, while SFT with $\mathcal{D}_{\text{UT}}$ provides substantial gains in best-of-$N$ improvement relative to the base Qwen3-4B, it yields only marginal improvements in unit test fidelity. We conjecture that this is because SFT with $\mathcal{D}_{\text{UT}}$ trains the LLM to generate unit tests without step-by-step reasoning, making it difficult for the model to produce logically structured test suites (i.e., ranging from basic test cases to highly discriminative edge cases), which are essential for achieving high unit test fidelity.
@@ 4.3 Superiority of UTRL over Supervised Learning Approaches
To investigate whether training LLMs with reasoning-annotated unit tests via SFT can improve unit test generation, we also evaluate SFT with $\mathcal{D}_{\text{reason+UT}}$ and compare it with UTRL. Figure 5 presents best-of-$N$ improvement (left) and unit test fidelity (right) for Qwen3-4B trained with SFT baselines versus the same model trained with UTRL. While SFT + $\mathcal{D}_{\text{reason+UT}}$ achieves an improved unit test fidelity score of 0.604, higher than the SFT + $\mathcal{D}_{\text{UT}}$, it falls behind 0.741 achieved by UTRL. Although SFT baselines directly optimize LLMs using unit test labels annotated by humans or more capable teacher models, they show limited generalization to evaluation tasks compared with UTRL. This observation is consistent with prior findings that SFT tends to memorize the training distribution, whereas RL promotes better generalization in reasoning tasks (Chu et al., 2025). Moreover, because collecting high-quality unit test annotations for SFT is costly, UTRL offers a more scalable and practical approach for training LLMs for unit test generation.
@@ 4.4 Effectiveness of UTRL in Training Code Generator
In UTRL, the code generator is trained jointly with the unit test generator to produce solutions that maximize the pass rate over the generated unit tests. To evaluate the effectiveness of UTRL for training the code generator, we compare Qwen3-4B trained with UTRL against the same model trained with RL using rewards defined as the pass rate over GT unit tests. Figure 8 shows code accuracy and code score averaged over 945 evaluation tasks across RL training steps. The code generator trained with UTRL achieves a code accuracy of 15.3% and a code score of 0.391, comparable to the 15.9% and 0.423 obtained when training with rewards from GT unit tests. These results indicate that UTRL is effective in training code generators, and that the unit tests it produces are sufficiently reliable to serve as rewards for code generation tasks.
@@ 4.5 Effect of Iterative training
Figure 7: Code score and accuracy on the TACO evaluation tasks. We evaluate Qwen3-4B code generator trained with UTRL (green) and compare it with the same model trained using rewards defined as the pass rate over GT unit tests (orange).
@@
We examine whether iterative training creates increasingly challenging discrimination tasks for the unit test generator, and whether the generator improves by learning from them. Figure 8 shows discrimination rewards averaged over evaluation tasks. In the first iteration, the unit test generator is trained against code produced by the initial code generator, and the reward saturates after about 50 steps. At this point, we stop training and update the code generator via RL to maximize its pass rate against the current unit tests. When training resumes in iteration 2, the reward drops from 0.626 to 0.375, indicating that the updated code generator now produces solutions that are harder to distinguish from ground truth, thereby creating a more challenging discrimination task for the unit test generator. As training continues in iteration 2, the unit test generator improves the reward from 0.375 to 0.447. Correspondingly, as shown in Figure 10, the unit test generator at iteration 2 produces unit tests that yield superior best-of-$N$ performance compared to iteration 1.

### v3

@@ Abstract
Unit testing is a core practice in programming, enabling systematic evaluation of programs produced by human developers or large language models (LLMs). Given the challenges in writing comprehensive unit tests, LLMs have been employed to automate unit test generation, yet methods for training LLMs to produce high-quality unit tests remain underexplored. In this work, we propose UTRL, a novel reinforcement learning (RL) framework that trains an LLM to generate high-quality unit test given a programming instruction. Our key idea is to iteratively train two LLMs, the unit test generator and the code generator, in an adversarial manner via RL: (1) the unit test generator is trained to maximize a discrimination reward, encouraging it to produce tests that reveal faults in the code generator’s solutions; and (2) the code generator is trained to maximize a code reward, encouraging it to produce solutions that pass the unit tests generated by the unit test generator. In our experiment, we demonstrate that unit tests generated by Qwen3-4B trained via UTRL show higher quality compared to unit tests generated by the same model trained via supervised fine-tuning on ground-truth unit tests, yielding code evaluations that more closely align with those induced by the ground-truth tests. Moreover, Qwen3-4B trained with UTRL outperforms frontier models like GPT-4.1 and GPT-4o in generating high-quality unit tests, highlighting the effectiveness of UTRL in training LLMs for the unit test generation. Code and relevant materials are available at our project page: https://dgjun32.github.io/UTRL.
@@ 2 Related Work
Unit testing has been extensively explored for reliable software development (Fraser & Arcuri, 2011; Alagarsamy et al., 2024). As LLMs show strong capability in programming, recent works have explored the automation of unit test generation using large language models (LLMs), with a focus on evaluating the test generation capability of LLMs (Jain et al., 2024a; Wang et al., 2024), iteratively refining the unit test by leveraging coverage analysis or mutation testing as a contextual feedback (Alshahwan et al., 2024; Altmayer Pizzorno & Berger, 2025; Dakhel et al., 2024; Foster et al., 2025), or utilizing co-evolutionary algorithm (Li et al., 2025). Another line of work investigates training LLMs for unit test generation via supervised learning, with a focus on constructing high-quality unit test annotations at scale. CodeRM (Ma et al., 2025) leverages stronger teacher models to generate unit tests automatically, while UTGEN (Prasad et al., 2025) proposes a data collection strategy that perturbs code solutions to create faulty variants, and then collects high-quality test cases that discriminate between the correct and faulty code. The work most closely related to ours is CURE (Wang et al., 2025a), which explores the co-evolution of code generation and test case generation capabilities via reinforcement learning. However, our approach differs from CURE in two key aspects: (1) Unlike CURE, which assumes access to a dataset annotated with ground-truth unit tests, UTRL requires only instruction-code pair dataset, which is readily available at large scale (Li et al., 2023; Lambert et al., 2024; Prasad et al., 2025), and (2) UTRL is a framework aimed at training LLMs to generate an optimal set of test cases, rather than a single test case.
@@
Adversarial learning and self-play
Adversarial learning provides a general framework in which models improve by competing against an adaptive opponent, even enabling model to learn complex data distribution (Goodfellow et al., 2014). Building on this idea, self-play has emerged as a powerful paradigm where the opponent is derived from the model itself. A line of works has demonstrated that symmetric self-play can yield superhuman performance in strategic decision-making (Silver et al., 2017), while another line of work explores asymmetric formulations in which agents generate and solve tasks to construct an automatic curriculum (OpenAI et al., 2021). With the rapid progress of LLMs across diverse domains, self-play has recently been extended to LLM-based systems, enabling improvement with minimal or even zero additional data. AZR introduce a zero-data self-play framework in which a task generator and a task solver co-evolve to improve the model’s reasoning ability (Zhao et al., 2025). In alignment, SPIN applies self-play to alignment by iteratively contrasting LLM-generated responses with human responses, enabling improvement without costly preference data (Chen et al., 2024). In code generation, Sol-Ver utilizes self-play between a solution generator and a test generator to jointly improve code and unit tests (Lin et al., 2025). In theorem proving, STP leverage self-play between a conjecture generator and a prover to enhance theorem-proving capability of LLM (Dong & Ma, 2025).
@@ 3.2 Training Unit Test Generator
Figure 3: Overview of the process of computing discrimination reward with respect to unit test $\mathcal{T}$. First, among the $5$ test cases in the unit test $\mathcal{T}$, test cases that pass under the ground-truth code $C^{*}$ (i.e., $T_{1},T_{3},T_{5}$) are filtered, forming a set of functionally valid test cases. Second, the discrimination reward is defined as a ratio of sampled code solutions that are detected by least one valid test case. In this figure, among 6 sampled code solutions, 4 code solutions ($C_{1},C_{3},C_{4},C_{5}$) do not pass at least one valid test case, resulting in the discrimination reward of 0.667 ($\frac{4}{6}$).
@@ 4.1 Experimental Setup
  • LLM baselines without fine-tuning: We use 4 open-source LLMs (Qwen3-4B, Qwen3-8B, Qwen3-14B, Qwen3-32B, Deepseek-Coder-V2-Lite; Qwen et al. 2025; Zhu et al. 2024) and 2 closed LLMs (GPT-4o, GPT-4.1; Hurst et al. 2024; OpenAI 2025) as baselines [^1] We note that 4 open-source LLMs used as baselines are instruction-finetuned models..
  • Supervised fine-tuning (SFT): We fine-tune Qwen3-4B with instruction–unit test pairs through supervised learning. For this purpose, we use high-quality unit tests from the TACO training dataset $\mathcal{D}_{\text{UT}}$. To investigate whether reasoning signals can improve unit test generation, we construct a reasoning-augmented dataset $\mathcal{D}_{\text{reason+UT}}$ containing unit tests paired with their corresponding reasoning. Specifically, we first generate reasoning and unit tests for each training task using Gemini-2.5-flash (Google Deepmind, 2025). [^2] Instead of attaching reasoning to the original unit tests in TACO, which led to frequent hallucinations and inconsistencies, we regenerate both reasoning and unit tests together. We then filter functionally valid reasoning–test case pairs and use them as target labels for SFT training.
@@
  • Ground-truth (GT) unit tests: We also compare against the unit tests provided in the benchamrk datasets (Li et al., 2023; Jain et al., 2024b), which we consider as ground truth (GT) and regard as an upper bound since they are rigorously verified and consist of a large number of high-quality test cases covering diverse edge cases.
@@
For evaluation, we utilize rigorously verified 945 competitive programming tasks provided in the TACO evaluation dataset (Li et al., 2023) and 511 programming tasks provided in the LiveCodeBench-v2 (Jain et al., 2024b). To measure the quality of the generated unit tests, we propose the following evaluation schemes:
  • Best-of-NN improvement: To examine whether the generated unit tests include high-quality test cases, we perform best-of-$N$ sampling for code generation, using the generated unit tests as evaluators. Specifically, we generate $N$ candidate solutions from code LLMs and select the one that passes the largest number of generated unit tests. To evaluate the quality of the selected code, we then measure two metrics: (1) code score, defined as the fraction of test cases in a GT unit test (i,e., high-quality unit tests from evaluation dataset) that the code solution passes, and (2) code accuracy, which indicates whether the solution passes all test cases in the GT unit test.
@@
We provide more details about the evaluation metrics in Appendix A.4.
4.2 Quality of Unit Tests
Best-of-NN improvement
Table 1 reports the best-of-$N$ improvement achieved by Qwen3-4B and Qwen3-14B trained with UTRL, compared against several baselines. When used as code evaluators for best-of-32 sampling with Qwen3-8B and Qwen3-14B code generators, unit tests produced by Qwen3-4B trained with UTRL yield code accuracies of 14.9% and 17.3%, respectively. For comparison, Qwen3-4B trained with SFT achieves only 11.7% and 14.0%. Moreover, Qwen3-4B trained with UTRL outperforms frontier proprietary models such as GPT-4.1 and GPT-4o, underscoring the effectiveness of UTRL in training LLMs to generate high-quality unit tests. With increased model capacity, Qwen3-14B trained with UTRL attains stronger performance, reaching code accuracies of 15.0% and 17.7%. Notably, UTRL is effective even when paired with closed-source code generators such as GPT-4o (see Table 8 in Appendix A.2 for supporting results).
@@ 4.2 Quality of Unit Tests
Qwen3-4B + UTRL (ours) 0.578 +0.232 14.9 +6.9 0.610 +0.208 17.3 +6.6
Qwen3-14B + UTRL (ours) 0.599 +0.253 15.0 +7.0 0.627 +0.225 17.7 +7.0
GT (upperbound) 0.668 +0.304 20.3 +13.1 0.699 +0.278 24.2 +13.7
@@
Figure 4 presents the fidelity of unit tests generated by Qwen3-4B trained with UTRL, compared against several baselines. Qwen3-14B and Qwen3-4B trained with UTRL achieves a Spearman’s correlation of 0.827 and 0.794, where Qwen3-14B trained via UTRL outperforms GPT-4.1, demonstrating the effectiveness of UTRL in training LLMs to generate unit tests that induce code evaluations consistent with the GT unit tests. Interestingly, while SFT with $\mathcal{D}_{\text{UT}}$ provides substantial gains in best-of-$N$ improvement relative to the base Qwen3-4B, it yields only marginal improvements in unit test fidelity. We conjecture that this is because SFT with $\mathcal{D}_{\text{UT}}$ trains the LLM to generate unit tests without step-by-step reasoning, making it difficult for the model to produce logically structured test suites (i.e., ranging from basic test cases to highly discriminative edge cases), which are essential for achieving high unit test fidelity.
@@ 4.3 Superiority of UTRL over Supervised Learning Approaches
To investigate whether training LLMs with reasoning-annotated unit tests via SFT can improve unit test generation, we also evaluate SFT with $\mathcal{D}_{\text{reason+UT}}$ and compare it with UTRL. Figure 5 presents best-of-$N$ improvement (left) and unit test fidelity (right) for Qwen3-4B trained with SFT baselines versus the same model trained with UTRL. While SFT + $\mathcal{D}_{\text{reason+UT}}$ achieves an improved unit test fidelity score of 0.604, higher than the SFT + $\mathcal{D}_{\text{UT}}$, it falls behind 0.741 achieved by UTRL. Additionally, in best-of-N improvement, SFT + $\mathcal{D}_{\text{reason+UT}}$ shows performance even inferior to SFT + $\mathcal{D}_{\text{UT}}$. We conjecture that this is because the training dataset $\mathcal{D}_{\text{reason+UT}}$ comprises synthetic test cases generated by Gemini-2.5-Flash rather than highly discriminative ground-truth unit tests that constitute $\mathcal{D}_{\text{UT}}$. Hence, the model trained with $\mathcal{D}_{\text{reason+UT}}$ generates relatively less discriminative unit tests compared to the same model trained with $\mathcal{D}_{\text{UT}}$, which results in inferior best-of-N improvement, since Best-of-N improvement is largely determined by the presence of highly discriminative test cases. Although SFT baselines directly optimize LLMs using unit test labels annotated by humans or more capable teacher models, they show limited generalization to evaluation tasks compared with UTRL. This observation is consistent with prior findings that SFT tends to memorize the training distribution, whereas RL promotes better generalization in reasoning intensive tasks (Chu et al., 2025). Moreover, since collecting high-quality unit test annotations for SFT is costly, UTRL offers a more practical approach for training LLMs for unit test generation.
@@ 4.4 Effectiveness of UTRL in Training Code Generator
In UTRL, the code generator is adversarially trained against the unit test generator to produce code solutions that maximize the pass rate over the generated unit tests. To evaluate the effectiveness of UTRL for training the code generator, we measure the pass@1 code accuracy of code solutions generated by Qwen3-4B trained via UTRL, and compare it against two code generator training baselines: (1) RL using rewards defined as the pass rate over unit tests generated by GPT-4o, and (2) SFT using ground-truth code solution provided in the training dataset. Moreover, we also evaluate Qwen3-4B trained via RL using rewards defined as the pass rate over GT unit tests provided in the TACO dataset, which we regard as our upperbound. Figure 8 shows pass@1 code accuracy averaged over 945 evaluation tasks. The code generator trained with UTRL achieves a code accuracy of 15.3%, which is remarkably higher than the two baselines, and even comparable to 15.9% achieved by the code generator trained to maximize pass rate over the GT unit tests. While training the code generator to maximize the pass rate over GPT-4o-generated unit tests improved the pass@1 accuracy to some extent, the code accuracy saturates under 12%, and we observe that directly fine tuning code generator with ground-truth code solutions via SFT severely degrades the performance on unseen evaluation tasks, resulting in merely 3.6% pass@1 code accuracy. These results indicate the effectiveness of UTRL in training the code generator, especially when high-quality ground-truth unit tests are infeasible.
@@ 4.5 Effect of Iterative training
Figure 7: Pass@1 accuracy of code generators on the TACO evaluation tasks. Qwen3-4B code generator trained with UTRL outperforms the same model trained with GPT-4o-generated unit tests and the same model trained via SFT in terms of pass@1 accuracy.
@@
We examine whether iterative training enables the code generator to produce near-perfect code solution, thereby providing increasingly challenging discrimination tasks for the unit test generator, and whether the generator improves by learning to discriminate the near-perfect code from ground-truth code. Figure 8 shows discrimination rewards averaged over evaluation tasks. In the first iteration, the unit test generator is trained against code produced by the initial code generator, and the discrimination reward saturates after about 50 steps. At this point, we stop training and update the code generator via RL to maximize its pass rate against the current unit tests. When training resumes in iteration 2, the discrimination reward drops from 0.626 to 0.375, indicating that the updated code generator now produces code solutions that are harder to distinguish from ground-truth, thereby creating a more challenging discrimination task for the unit test generator. As training continues in iteration 2, the unit test generator improves the reward from 0.375 to 0.447. Correspondingly, as shown in Figure 10, the unit test generator at iteration 2 produces unit tests that yield superior best-of-$N$ performance compared to iteration 1, even outperforming GPT-4.1. Furthermore, we extend the training to a third iteration, but we did not observe notable improvement in discrimination reward. While we conducted these experiments using Qwen3-4B as the unit test generator, we believe that employing a larger-capacity model as a unit test generator would likely allow the adversarial process to sustain meaningful improvements over more iterations.